### v4

@@ VI-B1 The effect of bit plane coding on grayscale image datasets
MNIST[44] 98.62 97.87 98.93 (+0.31)
KMNIST[55] 94.13 92.33 95.58 (+1.45)
Fashion-MNIST[56] 88.83 86.66 90.19 (+1.36)
@@
First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information. For that, we employed three popular datasets of grayscale images, including MNIST[44], KMNIST[55], Fashion-MNIST[56]. The results of MNIST dataset were already shown in Table III. However, they are still included here to see the behaviors of the coding methods over different datasets. As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel. This process would produce 8 bit-plane channels that represent the image information in binary form. In this part, we investigate three options,namely:
@@ VI-B2 Comparison of SNN variants
In addition, we also evaluated our spike coding scheme in different SNN variants. We employ only both grayscale and color image datasets in this part, including MNIST, KMNIST, Fashion-MNIST and CIFAR10. We benchmarked against multiple variants of Spiking ResNet[28] with LIF surrogate neurons, which have demonstrated competitive performance in prior work, along with AND and IAND variants of SEW-ResNet18[11].
In shown in Table V, the results reveal that in all of the tested variants, the proposed method consistently achieves a higher accuracy than the baseline. The performance gains range from 0.31% (in the case of SEW-ResNet18 (ADD) on MNIST dataset) to 44.40% (in the case of Spiking-ResNet50 on Fashion-MNIST). Also, Spiking-ResNet variants tend to underperform when compared with SEW-ResNet18 variants. The average accuracy values across all datasets show that SEW-ResNet18 (ADD) is the best variant. Fang et al. [11] also demonstrated that this variant usually achieves superior performance. Therefore, in subsequent experiments, we exclusively utilize this variant for benchmarking.
@@ VI-B3 The effect of bit plane coding on color images
CIFAR10[32] 70.69 37.28 73.49 (+2.80)
CIFAR100[32] 38.57 11.28 42.15 (+3.58)
Caltech101[57] 61.67 43.28 64.55 (+2.88)
Caltech256[58] 30.04 21.75 41.51 (+11.47)
EuroSAT[59] 84.39 79.87 88.48 (+4.09)
Imagenette[60] 72.99 48.20 78.68 (+5.69)
Food101[61] 19.62 6.29 36.95 (+17.33)
@@
For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[32], CIFAR100[32], EuroSAT[59], Caltech101[57], Caltech256[58], Imagenette[60] (which is a subset of ImageNet[62]) and Food101[61]. The default color model of these datasets is RGB. Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels. So, among these datasets, CIFAR10[32], CIFAR100[32] and EuroSAT[59] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
@@
CIFAR10[32] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
CIFAR100[32] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
Caltech101[57] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
Caltech256[58] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
EuroSAT[59] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
Imagenette[60] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
Food101[61] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
@@ VI-C Discussion
Our experimental results demonstrate that bit-plane coding can enhance the performance of SNNs. Despite the significant performance gains, the added computational complexity is a limitation of our method. As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some applications. Still, along with most spike coding methods utilized in surrogate training in SNN, both proposed method and baseline are still able to compute spike signal within linear time complexity $\mathcal{O}(n)$[3]. Empirical evaluation of processing speed and utilized memory could varied across datasets, attributed to factors such as input resolution, hardware specifications, and software library versions. Also, another limitation is that, similar to recent studies [28, 11], the experiments in our study are still based on software simulations. Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions. First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs for different machine learning tasks or cross ANN-SNN benchmarking[63]. Second, it is obvious that the choice of color model may depend on the task in question, such as temporal-based segmentation[64] , object recognition[65] and knowledge distillation [66]. So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks. Additionally for future works, the proposed method should be implemented by mapping proposed spike coding method on hardware platforms for deployment such as FPGA or Intel Loihi chip[6] and tested in practical scenarios such as low power edge devices[67].
@@ Appendix A Basis for stochastic theorem reusability in SNN surrogate training
Now, we consider an example theorem along with it’s auxiliary lemmas derived from a theoretical analysis of a gradient-based optimization experiment[45]:
@@
Since the components of Lemma A.2 do not depend on the distribution of $\mathbf{x}$, the lemma remains unchanged. However, Lemma A.3 no longer necessarily holds, as any multiplication involving $\mathbf{x}$ can be interpreted as a $d$-dimensional Littlewood–Offord problem[68], which only admits a probabilistic upper bound rather than a deterministic one. Consequently, Lemma A.3 becomes unstable and holds only under the additional assumption that the unit vectors also follow a standard Gaussian distribution, i.e., $\mathbf{w},\mathbf{v}\sim\mathcal{N}(0,1)$.
@@ Appendix B Proof for SNR inequality
Firstly, we will provide some background knowledge to assist our mathematical derivation of Equation 11. Given a SNN model that employ hard-reset IF neurons with backpropagate through time (BPTT)[11, 10, 48, 23, 22], optimization scheme is defined as:
@@
As noted from Appendix A, SNR expression denoted in Equation 10 [45] is adaptable to surrogate SNN model by substitute results from Equation 18 as:
@@
$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
@@
Under the property of expectation’s sum and Minkowski’s sum [69]:
@@
$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
@@
$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
@@
$g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim g_{\mathbf{w}}(\mathbf{x})$ (22)
@@
where each squared entry $\delta_{ij}^{2}$ would follow a chi-squared distribution with 1 degree of freedom with $\delta_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(1)$, the sum of all squared entries follows a chi-squared distribution with $n_{1}^{\delta}n_{2}^{\delta}$ degrees of freedom (by Cochran’s theorem[70]):
@@
Thus, $\left\|\delta({\mathbf{x}})\right\|_{2}$ grows asymptotically as $O(\left|\sigma_{\delta}\right|\sqrt{n_{1}^{\delta}n_{2}^{\delta}})$. We can also quantify the probability for lower bound and upper bound of $\left\|\delta({\mathbf{x}})\right\|$ using the consequences of the standard Laurent-Massart bounds[71]:

### v5

@@ VI-B1 The effect of bit plane coding on grayscale image datasets
MNIST[deng2012mnist] 98.62 97.87 98.93 (+0.31)
KMNIST[clanuwat2018deep] 94.13 92.33 95.58 (+1.45)
Fashion-MNIST[xiao2017fashion] 88.83 86.66 90.19 (+1.36)
@@
First, we aim to investigate whether the proposed approach yields any significant or insightful results without impacts of color information. For that, we employed three popular datasets of grayscale images, including MNIST[deng2012mnist], KMNIST[clanuwat2018deep], Fashion-MNIST[xiao2017fashion]. The results of MNIST dataset were already shown in Table III. However, they are still included here to see the behaviors of the coding methods over different datasets. As grayscale images consist of only intensity values ranging between [0, 255], we apply our coding process with one channel. This process would produce 8 bit-plane channels that represent the image information in binary form. In this part, we investigate three options,namely:
@@ VI-B2 Comparison of SNN variants
In addition, we also evaluated our spike coding scheme in different SNN variants. We employ only both grayscale and color image datasets in this part, including MNIST, KMNIST, Fashion-MNIST and CIFAR10. We benchmarked against multiple variants of Spiking ResNet[hu2021spiking] with LIF surrogate neurons, which have demonstrated competitive performance in prior work, along with AND and IAND variants of SEW-ResNet18[fang2021deep].
In shown in Table V, the results reveal that in all of the tested variants, the proposed method consistently achieves a higher accuracy than the baseline. The performance gains range from 0.31% (in the case of SEW-ResNet18 (ADD) on MNIST dataset) to 44.40% (in the case of Spiking-ResNet50 on Fashion-MNIST). Also, Spiking-ResNet variants tend to underperform when compared with SEW-ResNet18 variants when benchmarking under spike-based backpropagation (which had been priorly noted by [fang2021deep] due to gradient instability in previous works and Section II-B). The average accuracy values across all datasets show that SEW-ResNet18 (ADD) is the best variant. Fang et al. [fang2021deep] also demonstrated that this variant usually achieves superior performance. Therefore, in subsequent experiments, we exclusively utilize this variant for benchmarking.
@@ VI-B3 The effect of bit plane coding on color images
CIFAR10[alex2009learning] 70.69 37.28 73.49 (+2.80)
CIFAR100[alex2009learning] 38.57 11.28 42.15 (+3.58)
Caltech101[li_andreeto_ranzato_perona_2022] 61.67 43.28 64.55 (+2.88)
Caltech256[griffin_holub_perona_2022] 30.04 21.75 41.51 (+11.47)
EuroSAT[helber2019eurosat] 84.39 79.87 88.48 (+4.09)
Imagenette[imagenette] 72.99 48.20 78.68 (+5.69)
Food101[food101] 19.62 6.29 36.95 (+17.33)
@@
For our experiment, we conducted tests on various color-image classification datasets, including CIFAR10[alex2009learning], CIFAR100[alex2009learning], EuroSAT[helber2019eurosat], Caltech101[li_andreeto_ranzato_perona_2022], Caltech256[griffin_holub_perona_2022], Imagenette[imagenette] (which is a subset of ImageNet[deng2009imagenet]) and Food101[food101]. The default color model of these datasets is RGB. Due to computing resource constraints of the simulation tool, large images are downscaled to 224x224 pixels. So, among these datasets, CIFAR10[alex2009learning], CIFAR100[alex2009learning] and EuroSAT[helber2019eurosat] are trained with original input size while images of all other dataset are resized to 224x224 pixels.
@@
CIFAR10[alex2009learning] 73.49 73.73 73.64 73.39 73.06 73.63 74.06
CIFAR100[alex2009learning] 42.15 41.31 41.62 41.94 41.91 41.73 42.03
Caltech101[li_andreeto_ranzato_perona_2022] 64.55 65.65 65.53 66.46 68.41 65.30 66.22
Caltech256[griffin_holub_perona_2022] 41.51 40.65 40.84 40.55 39.94 40.50 40.79
EuroSAT[helber2019eurosat] 88.48 86.61 88.11 88.74 88.81 88.87 87.48
Imagenette[imagenette] 78.68 77.35 78.78 78.55 78.50 77.35 78.70
Food101[food101] 36.95 37.71 36.98 38.03 36.41 37.10 35.87
@@ VI-C Discussion
Our experimental results demonstrate that bit-plane coding can enhance the performance of SNNs. Despite the significant performance gains, the added computational complexity is a limitation of our method. As shown in Figure 5, the average complexity increase in terms of processing time can be from 48.57% to 85.35%, which could be a drawback in some edge applications where resource utilization are of higher priority. Still, along with most spike coding methods utilized in surrogate training in SNN, both proposed method and baseline are still able to compute spike signal within linear time complexity $\mathcal{O}(n)$[auge2021survey]. Empirical evaluation of processing speed and utilized memory could varied across datasets, attributed to factors such as input resolution, hardware specifications, and software library versions. Also, another limitation is that, similar to recent studies [hu2021spiking, fang2021deep], the experiments in our study are still based on software simulations. Implementations and experiments on hardware would be definitely needed to understand the actual complexity.
As our study is the first one to investigate bit plane coding in SNNs, it can be extended in different directions. First, a more extensive theoretical analysis is necessary to model the impact of bit plane coding on the performance of SNNs for different machine learning tasks or cross ANN-SNN benchmarking[deng2020rethinking]. Second, it is obvious that the choice of color model may depend on the task in question, such as temporal-based segmentation[patel2021spiking] , object recognition[su2023deep] and knowledge distillation [guo2023joint]. So, experiments in different application domains to understand the effects of the proposed method and color models for different tasks. Additionally for future works, the proposed method should be implemented by mapping proposed spike coding method on hardware platforms for deployment such as FPGA or Intel Loihi chip[nunes2022spiking] and tested in practical scenarios such as low power edge devices[zheng2021balancing].
@@ Appendix A Basis for stochastic theorem reusability in SNN surrogate training
Now, we consider an example theorem along with it’s auxiliary lemmas derived from a theoretical analysis of a gradient-based optimization experiment[shalev2017failures]:
@@
Since the components of Lemma A.2 do not depend on the distribution of $\mathbf{x}$, the lemma remains unchanged. However, Lemma A.3 no longer necessarily holds, as any multiplication involving $\mathbf{x}$ can be interpreted as a $d$-dimensional Littlewood–Offord problem[littlewood1939number], which only admits a probabilistic upper bound rather than a deterministic one. Consequently, Lemma A.3 becomes unstable and holds only under the additional assumption that the unit vectors also follow a standard Gaussian distribution, i.e., $\mathbf{w},\mathbf{v}\sim\mathcal{N}(0,1)$.
@@ Appendix B Proof for SNR inequality
Firstly, we will provide some background knowledge to assist our mathematical derivation of Equation 11. Given a SNN model that employ hard-reset IF neurons with backpropagate through time (BPTT)[fang2021deep, eshraghian2023training, doi:10.1126/sciadv.adi1480, luu2025hybrid, luu2025accuracy], optimization scheme is defined as:
@@
As noted from Appendix A, SNR expression denoted in Equation 10 [shalev2017failures] is adaptable to surrogate SNN model by substitute results from Equation 18 as:
@@
$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x}),$
@@
Under the property of expectation’s sum and Minkowski’s sum [zakon2004mathematical]:
@@
$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x}),$
@@
$h(\mathbf{x})\cdot g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g^{T}_{\mathbf{w}}(\mathbf{x}),$
@@
$g^{T+bit}_{\mathbf{w}}(\mathbf{x})\sim g^{T}_{\mathbf{w}}(\mathbf{x})$ (22)
@@
where each squared entry $\delta_{ij}^{2}$ would follow a chi-squared distribution with 1 degree of freedom with $\delta_{ij}^{2}\sim\sigma^{2}_{\delta}\chi^{2}(1)$, the sum of all squared entries follows a chi-squared distribution with $n_{1}^{\delta}n_{2}^{\delta}$ degrees of freedom (by Cochran’s theorem[cochran1934distribution]):
@@
Thus, $\left\|\delta({\mathbf{x}})\right\|_{2}$ grows asymptotically as $O(\left|\sigma_{\delta}\right|\sqrt{n_{1}^{\delta}n_{2}^{\delta}})$. We can also quantify the probability for lower bound and upper bound of $\left\|\delta({\mathbf{x}})\right\|$ using the consequences of the standard Laurent-Massart bounds[laurent2000adaptive]:
@@
[Uncaptioned image] Luu Trong Nhan (Member, IEEE) received the B.Sc. degree and M.Sc. degree in Computer Science from The University of Aizu, Japan, in 2024 and 2025 respectively. He is currently working as Research Assistant at College of Information and Communication Technology, Can Tho University, Vietnam. His research interests include machine learning, artificial intelligent algorithm, computational neuroscience and quantum computing.
[Uncaptioned image] Luu Trung Duong received the B.E. degree in Informatics from Can Tho University (CTU), Vietnam, in 1997 and M.Sc. degree in Mobile and Distributed Computer Network from Leeds Beckett University, England, in 2003. He has been working at Can Tho University since 1997 to present, starting as a Network Manager and currently is the Director of Center for Digital Transformation and Communication. His research interests include large scale networking solution design, computer security and network management.
[Uncaptioned image] Pham Ngoc Nam (Member, IEEE) received the bachelor’s degree in electronics engineering from the Hanoi University of Science and Technology (HUST), in 1997, and the M.S. degree in artificial intelligence and the Ph.D. degree in electrical engineering from KU Leuven, Belgium, in 1999 and 2004, respectively. He is currently the Vice Dean of the College of Engineering and Computer Science, VinUniversity, Vingroup. He is also a Visiting Scholar with Cornell University. He has been the PI of one key national project and three ministerial-level projects. He is the author or coauthor of 100 scientific articles, including more than 30 ISI and Scopus publications. His research interests include artificial intelligence, QoS/QoE management for multimedia applications, reconfigurable computing, and low-power embedded system design. He has been a key member of four other national projects.
[Uncaptioned image] Truong Cong Thang (Senior Member, IEEE) received the B.E. degree from Hanoi University of Technology, Vietnam, in 1997 and Ph.D. degree from KAIST, Korea, in 2006. From 1997 to 2000, he worked as an engineer in Vietnam Post & Telecom (VNPT). From 2006 to 2011, he was a Member of Research Staff at Electronics and Telecommunications Research Institute (ETRI), Korea. He was also an active member of Korean delegation to standard meetings of ISO/IEC and ITU-T from 2002 to 2011. Since 2011, he has been an Associate Professor of University of Aizu, Japan. His research interests include multimedia networking, image/video processing, content adaptation, IPTV and MPEG/ITU standards.